Q.2] Answer the following questions

Process

Burst Time

Priority

P1

3

1

P2

8

3

P3

2

4

P4

4

5

P5

5

1

  1. (21 points) Consider the set of processes shown in the table above, with the length of the CPU-burst time given in milliseconds. The processes are assumed to have arrived in the order P5, P4, P3, P2 , and P1, all approximately at time 0. Draw three Gantt charts illustrating the execution of these processes using SJF, Round-Robin (quantum size = 3), and a non-preemptive priority (a smaller priority number implies a higher priority) scheduling. Calculate average waiting time of each scheduling. Calculate average turnaround time of each scheduling.
  1. (1 points) What is convoy effect?
Answer #1

2) WT - TAT-BT Noless TOTCT-AT AI RT (1 3 5 22-3=14 22-0-22 22 8. O ܕܐ 2-2-0 2-0-2 Г. 2 O 2 9-4-5 P. 9-0 = 9 9 9 J IS 14-5 =91:) Round Robin ( Time Quantam =3 wT = TAT-BT. Process TAT = (T-AT Т AT BT CT 11 14 14 P. 14 22 22 22 8 6 8 8 8 P AM 2 13 P(it) non- preemptive Priority WT CT TAT Process Rienity BT 5 8 8 1 Pi 3 8 16 16 8 3 P₂ 16 18 18 2 4 wo 18 22 22 5 Pu 4 5 O 1if any doubts please comment.

please give like